# BAT: Learning Event-based Optical Flow with Bidirectional Adaptive Temporal Correlation

### v1

@@ 4.1 Experimental Setup
Figure 5: Future optical flow prediction. The first row illustrates events from timestamp $t_{i-1}$ to $t_{i}$, while the second row presents the optical flow results from $t_{i}$ to $t_{i+1}$. Given only past events, our method can predict accurate future optical flow.
4.2 Comparisons with State-of-the-art
@@
DSEC-Flow. Tab. 1 presents the comparison results on the DSEC-Flow benchmark. Our BAT ranks $1^{st}$ among all published methods for all metrics. Compared to other iterative methods, E-RAFT [6] and TMA [16], our BAT improves the 1PE metric by 39.45% and 28.98%, respectively. The 1PE metric is the most challenging evaluation metric, as it directly reflects the detailed performance of the predictions.
@@
Qualitative results on the DSEC-Flow benchmark are shown in Fig. 6. Due to the spatial sparsity of event data, previous methods [6, 16] tend to produce predictions with blurred edges and loss of details. In contrast, our BAT leverages bidirectional temporal motion cues, enabling it to predict sharp edges and preserve fine structures. Additionally, benefiting from the proposed backward temporal correlation, our BAT effectively handles occlusions caused by objects moving out of the target frame, as shown in Fig. 2.
@@
MVSEC. Following previous works, we perform training on outdoor_day2 sequence and evaluate on outdoor_day1 sequence. Tab. 2 presents the evaluation results corresponding to $dt=1$ grayscale frame and $dt=4$ grayscale frame. Our method achieves the highest prediction accuracy and outperforms other methods by a large margin. Specifically, for the more challenging $dt=4$ setting, our BAT surpasses the second-best method, TMA [16], by 34.26% for the %Out metric and 24.29% for the EPE metric.
@@ 4.3 Ablation Study
We conduct ablation studies to validate the effectiveness of the proposed components. All ablation models are trained on the DSEC-Flow training dataset and evaluated on the DSEC-Flow benchmark, as presented in Tab. 3. We use the TMA [16] with $N=3$ temporal groups as the Baseline. Baseline only computes forward temporal correlation, neglecting the backward motion cues. In contrast, the proposed bidirectional temporal correlation (BTC) computes both forward and backward temporal correlations, enabling the model to leverage bidirectional temporally dense motion cues for accurate and spatially detailed optical flow estimation. As a result, BTC achieves a significant improvement in accuracy. Furthermore, we introduce an adaptive temporal sampling (ATS) strategy to ensure temporal consistency when computing correlations, which enhances the model’s performance.
@@
Due to the presence of non-uniform motion, simply sampling the non-target event frames based on a linear motion assumption can introduce inconsistent temporal motion features. To mitigate this issue, we propose spatially adaptive temporal motion aggregation (SATMA), which efficiently and adaptively aggregates consistent target motion features while suppressing inconsistent motion features. As shown in Tab. 3, SATMA significantly improves the model’s prediction accuracy.
Attention type. To effectively perform temporal motion aggregation, we explore three types of attention, as shown in Tab. 4. Dense attention leads to excessive memory/computational costs and easily aggregates irrelevant temporal motion features that are beyond the regions of interest. To reduce memory/computational costs, we also explore spatial-reduction attention, which reduces the spatial scale of $\bm{K}$ and $\bm{V}$ through pooling before the attention. However, this causes a loss of detail in the target motion features, resulting in a significant degradation in accuracy. In comparison, deformable attention has low memory and computational demands while only aggregating relevant target motion features within the regions of interest.
@@
Sampling radius. Due to the rich temporal motion cues, even with using a sampling radius of 0, computing correlation at a single point, our method can still yield good results, as shown in Tab. 4. Better results are obtained with an increased radius.
@@
Iterations. By effectively integrating bidirectional and rich temporal motion cues, our results continuously improve with more iterations, as shown in Tab. 4 and Fig. 1.
4.4 Future Optical Flow Prediction
@@
Predicting future optical flow is a crucial task for interpreting and interacting with dynamic environments, such as assisting drones or autonomous vehicles in navigating through complex and dynamic environments. Previous methods [50, 6, 16] require the given $\mathcal{E}(t_{i},t_{i+1})$ to predict flow $\bm{f}_{t_{i}\rightarrow{t_{i+1}}}$, making them difficult to predict future optical flow. Benefiting from the proposed backward temporal correlation, our model can predict future optical flow $\bm{f}_{t_{i}\rightarrow{t_{i+1}}}$ using only the past event stream $\mathcal{E}(t_{i-1},t_{i})$, shown in Fig. 7. Our BAT (bwd corr) produces promising optical flow results on the DSEC-Flow benchmark, significantly outperforming the model-based method MultiCM [29], the self-supervised method TamingCM [26], and E-RAFT’s warm-start approach that initializes the flow for the next frame, as shown in Tab. 5. Visual results are shown in Fig. 5.
5 Conclusion

### v2

@@ Introduction
Figure 4: Future optical flow prediction. The first row illustrates events from timestamp $t_{i-1}$ to $t_{i}$, while the second row presents the optical flow results from $t_{i}$ to $t_{i+1}$. Given only past events, our method can predict future optical flow.
Comparisons with State-of-the-art
@@
DSEC-Flow. Table 1 presents the comparison results on the DSEC-Flow benchmark. Our BAT ranks $1^{st}$ among all published methods for all metrics. Compared to other iterative methods, E-RAFT (eraft) and TMA (tma), our BAT improves the 1PE metric by 39.45% and 28.98%, respectively. The 1PE metric is the most challenging evaluation metric, as it directly reflects the detailed performance of the predictions.
@@
Qualitative results on the DSEC-Flow benchmark are shown in Figure 5. Due to the spatial sparsity of event data, previous methods (eraft; tma) tend to produce predictions with blurred edges and loss of details. In contrast, our BAT leverages bidirectional temporal motion cues, enabling it to predict sharp edges and preserve fine structures. Additionally, benefiting from the proposed backward temporal correlation, our BAT effectively handles occlusions caused by objects moving out of the target frame, as shown in Figure 2.
@@
MVSEC. Following previous works, we perform training on outdoor_day2 sequence and evaluate on outdoor_day1 sequence. Table 2 presents the evaluation results corresponding to $dt=1$ grayscale frame and $dt=4$ grayscale frame. Our method achieves the highest prediction accuracy and outperforms other methods by a large margin. Specifically, for the more challenging $dt=4$ setting, our BAT surpasses the second-best method, TMA (tma), by 34.26% for the %Out metric and 24.29% for the EPE metric.
@@
We conduct ablation studies to validate the effectiveness of the proposed components. All ablation models are trained on the DSEC-Flow training dataset and evaluated on the DSEC-Flow benchmark, as presented in Table 3. We use the TMA (tma) with $N=3$ temporal groups as the Baseline. Baseline only computes forward temporal correlation, neglecting the backward motion cues. In contrast, the proposed bidirectional temporal correlation (BTC) computes both forward and backward temporal correlations, enabling the model to leverage bidirectional temporally dense motion cues for accurate and spatially detailed optical flow estimation. As a result, BTC achieves a significant improvement in accuracy. Furthermore, we introduce an adaptive temporal sampling (ATS) strategy to ensure temporal consistency when computing correlations, which enhances the model’s performance.
@@
Due to the presence of non-uniform motion, simply sampling the non-target event frames based on a linear motion assumption can introduce inconsistent temporal motion features. To mitigate this issue, we propose spatially adaptive temporal motion aggregation (SATMA), which efficiently and adaptively aggregates consistent target motion features while suppressing inconsistent motion features. As shown in Table 3, SATMA significantly improves the model’s prediction accuracy.
Attention type. To perform temporal motion aggregation effectively, we compare three attention types (Table 4). Dense attention incurs high memory/computation and introduces irrelevant features. Spatial-reduction attention lowers the cost by pooling $\bm{K}$ and $\bm{V}$, but sacrifices detail and degrades accuracy. In contrast, deformable attention efficiently focuses on relevant motion features with low overhead.
@@
Sampling radius. Due to the rich temporal motion cues, even with using a sampling radius of 0, computing correlation at a single point, our method can still yield good results, as shown in Table 4. Better results are obtained with an increased radius.
@@
Iterations. By effectively integrating bidirectional and rich temporal motion cues, our results continuously improve with more iterations, as shown in Table 4 and Figure 1.
Future Optical Flow Prediction
@@
Predicting future optical flow is a crucial task for interpreting and interacting with dynamic environments, such as assisting drones or autonomous vehicles in navigating through complex and dynamic environments. Previous methods (evflownet; eraft; tma) require the given $\mathcal{E}(t_{i},t_{i+1})$ to predict flow $\bm{f}_{t_{i}\rightarrow{t_{i+1}}}$, making them difficult to predict future optical flow. Benefiting from the proposed backward temporal correlation, our model can predict future optical flow $\bm{f}_{t_{i}\rightarrow{t_{i+1}}}$ using only the past event stream $\mathcal{E}(t_{i-1},t_{i})$, shown in Figure 6. Our BAT (bwd corr) produces promising optical flow results on the DSEC-Flow benchmark, significantly outperforming the model-based method MultiCM, the self-supervised method TamingCM, and E-RAFT’s warm-start approach that initializes the flow for the next frame, as shown in Table 5. Visual results are shown in Figure 4.
Conclusion
@@
Acknowledgments
This research is supported by the National Key R&D Program of China (2024YFE0217700), the National Natural Science Foundation of China (623B2036,62472184), the Fundamental Research Funds for the Central Universities, and the Innovation Project of Optics Valley Laboratory (Grant No. OVL2025YZ005).
Figure 7: More qualitative results on DSEC-Flow (dsec). Significant improvements are highlighted by red boxes. Images are provided for visualization only, since the optical flow is event-based. Our BAT effectively captures subtle details, while TMA (tma) struggles to predict them accurately.
Figure 8: Qualitative comparisons with TMA on MVSEC.
Figure 9: Spatially Adaptive Temporal Motion Aggregation (SATMA).
Details of Iterative Update
After performing the spatially adaptive temporal motion aggregation, we obtain forward fused motion features and backward fused motion features. These motion features are then concatenated to form the overall bidirectional motion features, represented as:
$\bm{M}^{bid}=\text{Concat}(\bm{M}^{fwd},\bm{M}^{bwd}),$ (10)
where $\bm{M}^{fwd}$ and $\bm{M}^{bwd}$ denote the forward and backward fused motion features, respectively, $\bm{M}^{bid}$ denotes bidirectional overall motion features. $\text{Concat}(\cdot,\cdot)$ represents the concatenation operation.
In addition to the bidirectional temporal motion features $\bm{M}^{bid}$, we also use a context network to extract context features $\bm{C}_{F}$. The context network takes event frames from timestamp $t_{i}-\Delta t/N$ to $t_{i+1}$ as input, and its architecture is identical to that of the feature extraction network. We then concatenate the $\bm{M}^{bid}$ and $\bm{C}_{F}$ to form $x_{k}$. Finally, the iterative update is performed using two ConvGRUs: one with a $1\times 5$ convolution and one with a $5\times 1$ convolution,
$z_{k}=$ $\;\sigma(\text{Conv}_{1\times 5}([h_{k-1},x_{k}],W_{z})),$ (11)
$r_{k}=$ $\;\sigma(\text{Conv}_{1\times 5}([h_{k-1},x_{k}],W_{r})),$
$\tilde{h}_{k}=$ $\,\tanh(\text{Conv}_{1\times 5}([r_{k}\odot h_{k-1},x_{k}],W_{h})),$
$h_{k}=$ $\;(1-z_{k})\odot h_{k-1}+z_{k}\odot\tilde{h}_{k},$
$z_{k}=$ $\;\sigma(\text{Conv}_{5\times 1}([h_{k},x_{k}],W_{z})),$
$r_{k}=$ $\;\sigma(\text{Conv}_{5\times 1}([h_{k},x_{k}],W_{r})),$
$\tilde{h}_{k}=$ $\,\tanh(\text{Conv}_{5\times 1}([r_{k}\odot h_{k},x_{k}],W_{h})),$
$h_{k}=$ $\;(1-z_{k})\odot h_{k}+z_{k}\odot\tilde{h}_{k},$
where $\sigma$ denotes the sigmoidsigmoid function, $W_{z}$, $W_{r}$, and $W_{h}$ are the parameters of the network. Based on the hidden state $h_{k}$, we decode a residual optical flow $\Delta\bm{f}$ through two convolutional layers. The $\Delta\bm{f}$ is added to the current optical flow estimation to update the optical flow.
More Visual Results
More comparisons with previous methods
As shown in Figure 7, we present additional qualitative results to further demonstrate the superiority of BAT. Compared with TMA (tma), our BAT captures subtle details more effectively.
Qualitative comparisons on MVSEC
We provide visual comparisons with TMA (tma) on the MVSEC (mvsec) dataset. Our method produces high-quality flow results while preserving fine structures, as shown in Figure 8.